support a person is selling goods to two different groups of people and the demand function for the first group is given as q1=18000-100p1. the demand function for the second group is q2=17000-100p2. the marginal cost of each group is $5.00. what is the profit if they cannot successfully segment the market for the two group of people? what will be the profit if taken successfully discriminate against 2 group of people? 


1
Expert's answer
2021-09-15T11:28:27-0400

If they cannot successfully segment the market for the two group of people, then:

p1 = p2 and

Q = q1 + q2 = 18000 - 100p + 17000 - 100p = 35000 - 200p,

p = 175 - 0.005Q,

MR = TR'(Q) = 175 - 0.01Q,

MR = MC,

175 - 0.01Q = 5,

Q = 17500 units,

p = 175 - 0.005×17500 = 87.5.

TP = (87.5 - 5)×17500 = 1,443,750.

If they successfully discriminate against 2 group of people, then:

p1 = 180 - 0.01q1,

MR1 = 180 - 0.02q1,

180 - 0.02q1 = 5,

0.02q1 = 175,

q1 = 8750,

p1 = 180 - 0.01×8750 = 92.5.

p2 = 170 - 0.01q2,

MR2 = 170 - 0.02q2,

170 - 0.02q2 = 5,

0.02q2 = 165,

q2 = 8250,

p2 = 170 - 0.01×8250 = 87.5.

TP = (92.5 - 5)×8750 + (87.5 - 5)×8250 = 1,446,250.

So, the total profit is higher, if they can discriminate.